OpenStudy (anonymous):

what is 1/4th of 40 feet?

OpenStudy (anonymous):

what is 1/4th of 30 feet i mean

OpenStudy (anonymous):

1/4th of 40 would be 10

OpenStudy (ash2326):

1/4th of 30 feet \[ \frac{1}{4} \times 30= \frac{30}{4}=7.5 feet\]

OpenStudy (anonymous):

7.5 is right :)

OpenStudy (anonymous):

thanks